# Rapidly-rotating Population III stellar models as a source of primary nitrogen

### v1

@@ 6 Discussion
In this section, we shall discuss three points: How the primary nitrogen production depends on the metallicity considering the whole range from $Z=0$ to $Z=0.020$; How the present yields compare with other works; How our yields impact the variation of N/O as a function of O/H when incorporated into a simple chemical evolution model.
@@ 6.1 The metallicity dependence
Figure 11: The stellar yields of nitrogen for the 9 (blue) and 20 $M_{\odot}$ (orange) models at different metallicities ($0\leqslant Z\leqslant 0.020$). The circles connected by the solid curve correspond to models with an initial rotation of $\upsilon_{\rm ini}=0.4\,\upsilon_{\rm crit}$ and the diamonds of $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$. The moderately-rotating models are from Murphy et al. (2021a, $Z=0$), Sibony et al. (submitted, $Z=10^{-5}$), Groh et al. (2019, $Z=0.0004$), Georgy et al. (2013a, $Z=0.002$), Ekström et al. (2012, $Z=0.014$), Eggenberger et al. (2021, $Z=0.006$), and Yusof et al. (2022, $Z=0.020$).
@@
Figure 12: Evolution of N/O as a function of O/H ratios for stellar models in the metallicity range of $0\leqslant Z\leqslant 0.020$ (coloured curves) and observed abundances of the Milky Way’s stars (grey points). Yields from models with different initial rotations are used (see Sect. 6.3). The observed data have been taken from the SAGA database (Suda et al., 2008, 2011; Yamada et al., 2013). The yellow star represent the Sun’s abundances from Asplund et al. (2009).
@@
Figure 11 shows how the stellar yields vary as a function of the initial metallicity for the 9 (orange) and 20 $M_{\odot}$ (blue) models. The solid curves show the results when the initial rotation is $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ at every metallicity, while the orange diamonds indicate the results when fast rotation is considered for $Z=0$.
@@
Let us first discuss the case of the 20 $M_{\odot}$ models. Schematically, we see that for the metallicity range between about 0.0004 and 0.020 with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$, the yields in nitrogen remain modest. They are even negative in the upper metallicity range. For those metallicities, nitrogen is mainly produced by the transformation of the initial carbon and oxygen abundances and nearly no nitrogen is produced by the diffusion of carbon and oxygen produced in the He-core into the H-burning shell. Yields are negative at high $Z$ because the mass of the ejecta where nitrogen is destroyed (typically matter processed by He-burning) is larger than the mass of the ejecta where nitrogen is produced.
@@
It is only in the metallicity range between 0 and 0.0004 that primary nitrogen is produced thanks to rotational mixing. Starting from $Z=0$, the increase of the yield with $Z$ is mainly due to rotational mixing that is favoured in slightly metal-rich models. As explained before, the differential rotation that drives the mixing increases when the metallicity increases. Beyond the peak, the decrease of the yield when the metallicity increases further is due to the larger and larger distance between the He- and H-burning zones. The 20 $M_{\odot}$ rapidly-rotating model is available only for $Z=0$. For that model, the primary nitrogen production reaches a level that is larger than the peak value shown at a metallicity $Z=10^{-5}$ with a moderate initial rotation.
@@
The 9 $M_{\odot}$ Pop III models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ do not show any strong peak between $Z=0$ and 0.0004. Primary nitrogen production in this metallicity range depends sensitively on the initial mass. It is interesting to note here that the central temperatures during the H-burning phase are smaller in less massive stars. Thus, decreasing the mass, everything else being kept constant, also tends to increase the change of entropy between the He- and H-burning zones making the mixing more difficult between the two zones.
@@ 6.2 Comparisons with other works
Comparing the nitrogen yields in Fig. 9 for the non-rotating models computed with Genec by Murphy et al. (2021a) (dashed blue curve) with those of Limongi & Chieffi (2012) (blue dot-dashed curve), one can note the following points. First, our yields for the 9 $M_{\odot}$ model are in line with those of Limongi & Chieffi (2012). Qualitatively we also obtain a bump that begins around 20 $M_{\odot}$. Since our grid does not consider the same initial masses, it is difficult to say more at the moment. We can just add that our yield for the 85 $M_{\odot}$ is also in line with the yield given by Limongi & Chieffi (2012).
If we compare the nitrogen yields by Ekström et al. (2010) for rotating models computed with Genec with those presented in this work, we obtain significant differences. This is in part due to the faster initial rotation considered here but also due to the fact that we used in the present models different diffusion coefficients for rotational mixing.
The differences between the yields in oxygen are mainly due to the different methods used for estimating the remnant mass. It is interesting to note that yields of Ekström et al. (2008) for rotating models and those of Limongi & Chieffi (2012) for non-rotating Pop III stars are very similar. Actually, in these two papers, the authors used remnant masses that are in general much smaller than those in this work. This gives larger oxygen yields that do not differ much between the rotating and non-rotating models.
These yields differ significantly from those obtained with the rapidly-rotating models computed here and even more from the non-rotating models of Murphy et al. (2021a). As explained already, this comes from the use of the method by Patton & Sukhbold (2020) for estimating the remnant mass.
@@ 6.3 Impact on the early chemical evolution of galaxies
Figure 12 shows the impact of the yields with different initial rotations in a simple galactic chemical evolution (GCE) model. Here, we computed the GCE with the one-zone closed-box model compiled in the chemical evolution library (celib, Saitoh, 2017; Hirai et al., 2021). This model adopts the initial mass function of Kroupa (2001) for 0.1--120 $M_{\odot}$ and stellar lifetimes in Portinari et al. (1998). We assume stars are constantly formed for 13 Gyr from a gas cloud of $10^{11}$ $M_{\odot}$ with a star formation efficiency of 0.08 Gyr-1. Note that rapidly-rotating models ($\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$) have been computed only for Pop III models. Therefore, we adopt the rapidly-rotating model for $Z$ = 0 and the rotating model with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ from Ekström et al. (2012); Georgy et al. (2013a); Groh et al. (2019); Eggenberger et al. (2021); Yusof et al. (2022); Sibony et al. (submitted) for metallicities above zero in the green curve. The red curve corresponds to the models above and to the models from Murphy et al. (2021a) with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ for the metallicity $Z=0$. Lastly, the blue curve corresponds to non-rotating models from all the works above. Observed data of the Milky Way’s stars were taken from the SAGA database (Suda et al., 2008, 2011; Yamada et al., 2013). We excluded CEMP stars with [C/Fe] >+absent>+> +0.7, which might be affected by binary mass transfer.
@@ 7 Conclusion
We have explored the impact of fast rotation on the evolution of Pop III stars with some discussion on their chemical feedback. The main conclusions are briefly synthesised below:
@@
    The appearance of the extended intermediate convective zone induced by the injection of carbon and oxygen produced in the He-burning region into the H-burning shell by diffusion/convection may have a dramatic impact on the structure and the evolution of the star. This intermediate convective zone may quench the He-burning in some model in the core for a while or at least significantly reduce the He-burning core (a process called CNO shell boost, as in Ekström et al., 2008). This has an impact on the nature of the stellar remnant and on nucleosynthesis.
@@
    Models with a mass equal to or larger than 60 $M_{\odot}$ spend more than half of their MS lifetime at the critical velocity limit. However, the mass lost mechanically remains very modest, the most massive model (120 $M_{\odot}$) loses 2% of its initial mass at the end of its evolution.
@@
    With the physics used here, models starting with 70% of the critical velocity on the ZAMS show no evidence of following a chemically homogeneous evolution trend. Actually, their tracks during the MS phase follow the classical redward evolution with the models rotating faster showing larger extension towards lower effective temperatures than slower rotators. The impact of fast rotation on the ionisation outputs appears modest in the present models.
@@
    Depending on the initial rotation at the ZAMS, a Pop III star can end its stellar lifetime as a BSG or a RSG. The rapidly-rotating models will be RSGs, as opposed to the BSGs from the slower-rotating models (for $M_{ini}>20$ $M_{\odot}$).
@@
This work illustrates the fact that the physics of rotational mixing is a key point to resolve in order to predict stellar yields, especially for the first generations of stars. The physics of rotation is in our view as important as the impact of close binary evolution. Indeed, if mass transfer in close binaries may strongly impact the evolution of stars, primary nitrogen production is essentially a question related to the internal mixing. Rotation appears as an excellent candidate for driving this mixing (note that close binary evolution and rotation are tightly intertwined through the exchange of angular momentum between orbits and spins of stars). An interesting feature of the rotational mixing is that it shows a dependence on the metallicity quite in line with the fact that important sources of primary nitrogen are needed only at low metallicities (see e.g. the discussion in Chiappini et al., 2005b).
@@
This work also underlines the importance of the initial rotation of Pop III stars. Moreover, even if we had a good knowledge of the rotational velocity distribution in the first Pop III stars, significant uncertainties persist regarding the physics of rotation. Similar work as discussed here should be done with other approaches for describing the transport of angular momentum and of chemical species by rotation and also by other types of instabilities (improvement in the physics of convection, internal waves, magnetic instabilities). There is some hope that testing stellar physics on stars that we can observe in the present-day Universe will allow us to constrain these mechanisms in a still tighter way. Then one can assume that this physics will also apply to Pop III stellar models. At the moment and likely for a still very long period there will be no way of directly measuring the rotation periods of Pop III stars from observations. Thus the only way to get information about that is to observe the consequences of stellar models computed with various initial rotations, ideally with different physics for the transport mechanisms, on the early chemical evolution of galaxies (as the origin of the chemical composition of CEMP stars and of the bulk of the galactic halo stars). A possible exception could be provided by cases of extreme flux amplification due to strong gravitational lensing (e.g. Zackrisson et al., 2015). Recently, such extreme lensing has been invoked to explain the Earendel stellar source (Welch et al., 2022), which has a non-negligible probability of being Pop III, although a metal-poor, massive Pop II star is more likely (Schauer et al., 2022). Other channels for getting information on the evolution of the first massive star generations is through the predictions of the rate of various types of supernovae (see the reviews by Frebel & Norris, 2015; Meynet & Maeder, 2017, and references therein). Recently nitrogen-rich high redshift galaxies have been observed (Oesch et al., 2016). Implications of the present rapidly-rotating models in this context are discussed in another paper (Nandal et al. in prep.).
@@ Appendix A Additional figures
A.1 Case of the rapidly-rotating 120 $M_{\odot}$ model
Figure 13 compares the variation of $U_{\rm r}$, $\Omega$, $D_{\rm shear}$ and $D_{\rm eff}$ as a function of the Lagrangian mass during the core H-burning phase when the mass fraction of hydrogen $X_{\rm c}$ is equal to 0.35 for the 120 $M_{\odot}$ models with a moderate (red curves) and a fast (green curves) initial rotation. The rapidly-rotating model shows stronger meridional currents (see the green curve in the upper left panel), which transport angular momentum from inside to outside. The value of $D_{\rm eff}$ is larger by many orders of magnitudes than $D_{\rm shear}$ in the radiative envelope and governs the mixing of the elements in that region.
Figure 13: Variation of $U_{\rm r}$ (upper left panel), $\Omega$ (upper right panel), $D_{\rm shear}$ (solid curves) and $D_{\rm eff}$ (dashed curves) (lower panel) as a function of the Lagrangian mass coordinate in the Pop III 120 $M_{\odot}$ models with different initial rotations. The moderately-rotating model was taken from Murphy et al. (2021a).
@@ A.2 CNO shell boost in a Pop III 60 $M_{\odot}$ model
Figure 14: Comparing stages for the 60 $M_{\odot}$ rapidly-rotating model as in Fig. 4.
The upper panel of Fig. 14 shows the HRD of the rapidly-rotating Pop III 60 $M_{\odot}$ model. Coloured points are overplotted. The orange point indicates when the core He-burning phase begins, while the purple point indicated the stage when an extended intermediate convective zone appears as a result of injections of carbon and oxygen from the He-burning core to the H-burning shell. Meanwhile, the convective He-burning core is reduced. Those points correspond to the vertical coloured lines in the lower Kippenhahn plot, respectively.
@@
In Fig. 15, we have an example of 20 $M_{\odot}$ Pop III models with zero and moderate rotation. We notice that the H- and He-burning shells are very close to each other in the non-rotating model. More precisely, in the non-rotating model, the mass separating the H- to the He-burning shell is a fraction of 1 $M_{\odot}$ during the core C-burning phase. While, in the moderately-rotating model, during the same phase, the mass separating the H- and He-burning shell is more than 2 $M_{\odot}$. This is in our view an indirect consequence of the fact that in the moderately-rotating model (right panel), a small convective zone occurs above the H-shell at the very beginning of the core He-burning phase (due to the rotational mixing some C and O migrate into the H-shell and boost its energy production). This has for an effect to make the mass of the convective He-core smaller during the first part of the He-burning phase. If the H-shell is replenished in hydrogen by the convective shell, it stays longer in deeper regions of the star, so it migrates slowly towards outer regions. In the non-rotating model, the H-burning shell at the beginning of the core He-burning phase migrates much faster outwards than in the rotating model (the shell is not replenished by the intermediate convective zone and thus progresses more rapidly outwards). This makes the He-core to also increases more rapidly in mass. As a result of this behaviour the two shells, just after the end of the core He-burning phase are separated by less mass in the non-rotating model than in the rotating one. Compare, for instance, in the two Kippenhahn diagram (see Fig. 15) the mass between the blue and green solid lines at a time 4. Due to the greater proximity in the non-rotating model, a CNO shell boost occurs. As a result, we see that the H-shell shifts to a position at a smaller lagrangian mass coordinate. So we see how complex can be this behaviour. In the rotating model a modest early CNO boost occurs that create conditions that disfavour any further CNO shell boost. In the rotating model the early CNO shell boost does not occur in absence of any rotational mixing, but it creates conditions more favourable for a CNO boost occurring at the end of the core He-burning phase.
Moreover, at the end of core He-burning phase, the moderately-rotating mode has higher nitrogen yield (approximately the moderately-rotating has ${}^{14}{\rm N}=3.2\times 10^{-7}$, while the non-rotating has ${}^{14}{\rm N}=1.2\times 10^{-7}$). Therefore, we see that those yields on the Fig. 9 are a result after the core He-burning phase.

### v2

@@ 6 Discussion
In this section, we discuss three points: how the primary nitrogen production depends on the metallicity, considering the whole range from $Z=0$ to $Z=0.020$; how the present yields compare with other works; and how our yields impact the variation of N/O as a function of O/H when incorporated into a simple chemical evolution model.
@@ 6.1 The metallicity dependence
Figure 11: Stellar yields of nitrogen for the 9 (blue) and 20 $M_{\odot}$ (orange) models at different metallicities ($0\leqslant Z\leqslant 0.020$). The circles connected by the solid curve correspond to models with an initial rotation of $\upsilon_{\rm ini}=0.4\,\upsilon_{\rm crit}$ and the diamonds ones with $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$. The moderately rotating models are from Murphy et al. (2021a, $Z=0$), Sibony et al. (2014, $Z=10^{-5}$), Groh et al. (2019, $Z=0.0004$), Georgy et al. (2013a, $Z=0.002$), Ekström et al. (2012, $Z=0.014$), Eggenberger et al. (2021, $Z=0.006$), and Yusof et al. (2022, $Z=0.020$).
@@
Figure 12: Evolution of N/O as a function of O/H ratios for stellar models in the metallicity range of $0\leqslant Z\leqslant 0.020$ (coloured curves) and observed abundances of the Milky Way’s stars (grey points). Yields from models with different initial rotations were used (see Sect. 6.3). The observed data were taken from the SAGA database (Suda et al., 2008, 2011; Yamada et al., 2013). The yellow star represent the Sun’s abundances from Asplund et al. (2009).
@@
Figure 11 shows how the stellar yields vary as a function of the initial metallicity for the 9 (orange) and 20 $M_{\odot}$ (blue) models. The solid curves show the results when the initial rotation is $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ at every metallicity, while the diamonds indicate the results when fast rotation is considered for $Z=0$.
@@
Let us first discuss the case of the 20 $M_{\odot}$ models. Schematically, we see that for the metallicity range between about 0.0004 and 0.020 with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$, the yields in nitrogen remain modest. They are even negative in the upper metallicity range. For those metallicities, nitrogen is mainly produced by the transformation of the initial carbon and oxygen abundances and nearly no nitrogen is produced by the diffusion of carbon and oxygen produced in the He-core into the H-burning shell. Yields are negative at high $Z$ because the mass of the ejecta where nitrogen is destroyed (typically, matter processed by He-burning) is larger than the mass of the ejecta where nitrogen is produced.
@@
It is only in the metallicity range between 0 and 0.0004 that primary nitrogen is produced thanks to rotational mixing. Starting from $Z=0$, the increase in the yield with $Z$ is mainly due to rotational mixing, which is favoured in slightly metal-rich models. As was explained before, the differential rotation that drives the mixing increases when the metallicity increases. Beyond the peak, the decrease in the yield when the metallicity increases further is due to the larger and larger distance between the He- and H-burning zones. The 20 $M_{\odot}$ rapidly rotating model is available only for $Z=0$. For that model, the primary nitrogen production reaches a level that is larger than the peak value shown at a metallicity of $Z=10^{-5}$ with a moderate initial rotation.
@@
The 9 $M_{\odot}$ Pop III models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ do not show any strong peak between $Z=0$ and 0.0004. Primary nitrogen production in this metallicity range depends sensitively on the initial mass. It is interesting to note here that the central temperatures during the H-burning phase are smaller in less massive stars. Thus, decreasing the mass, everything else being kept constant, also tends to increase the change in entropy between the He- and H-burning zones, making mixing more difficult between the two zones.
@@ 6.2 Comparisons with other works
Comparing the nitrogen yields in Fig. 9 for the non-rotating models computed with Genec by Murphy et al. (2021a) (dashed blue curve) with those of Limongi & Chieffi (2012) (dash-dotted blue curve), one can note the following points. First, our yields for the 9 $M_{\odot}$ model are in line with those of Limongi & Chieffi (2012). Qualitatively, we also obtain a bump that begins around 20 $M_{\odot}$. Since our grid does not consider the same initial masses, it is difficult to say more at the moment. We can just add that our yield for the 85 $M_{\odot}$ is also in line with the yield given by Limongi & Chieffi (2012).
If we compare the nitrogen yields by Ekström et al. (2010) for rotating models computed with Genec with those presented in this work, we obtain significant differences. This is in part due to the faster initial rotation considered here but also due to the fact that in the present models we used different diffusion coefficients for rotational mixing.
The differences between the yields in oxygen are mainly due to the different methods used to estimate the remnant mass. It is interesting to note that yields of Ekström et al. (2008) for rotating models and those of Limongi & Chieffi (2012) for non-rotating Pop III stars are very similar. Actually, in these two papers, the authors used remnant masses that are in general much smaller than those in this work. This gives larger oxygen yields that do not differ much between the rotating and non-rotating models.
These yields differ significantly from those obtained with the rapidly rotating models computed here and even more from the non-rotating models of Murphy et al. (2021a). As was explained already, this comes from the use of the method of Patton & Sukhbold (2020) to estimate the remnant mass.
@@ 6.3 Impact on the early chemical evolution of galaxies
Figure 12 shows the impact of the yields with different initial rotations in a simple galactic chemical evolution (GCE) model. Here, we computed the GCE with the one-zone closed-box model compiled in the chemical evolution library (celib, Saitoh, 2017; Hirai et al., 2021). This model adopts the initial mass function of Kroupa (2001) for 0.1--120 $M_{\odot}$ and stellar lifetimes in Portinari et al. (1998). We assume stars are constantly formed for 13 Gyr from a gas cloud of $10^{11}$ $M_{\odot}$ with a star formation efficiency of 0.08 Gyr-1. We note that rapidly rotating models ($\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$) were computed only for Pop III models. Therefore, we adopt the rapidly rotating model for $Z$ = 0 and the rotating model with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ from Ekström et al. (2012); Georgy et al. (2013a); Groh et al. (2019); Eggenberger et al. (2021); Yusof et al. (2022); Sibony et al. (2024) for metallicities above zero in the green curve. The red curve corresponds to the models above and to the models from Murphy et al. (2021a) with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ for the metallicity $Z=0$. Lastly, the blue curve corresponds to non-rotating models from all the works above. Observed data of the Milky Way’s stars were taken from the SAGA database (Suda et al., 2008, 2011; Yamada et al., 2013). We excluded CEMP stars with [C/Fe] >+absent>+> +0.7, which might be affected by binary mass transfer.
@@ 7 Conclusion
We have explored the impact of fast rotation on the evolution of Pop III stars with some discussion of their chemical feedback. The main conclusions are briefly synthesised below:
@@
    The appearance of the extended intermediate convective zone induced by the injection of carbon and oxygen produced in the He-burning region into the H-burning shell by diffusion or convection may have a dramatic impact on the structure and the evolution of the star. This intermediate convective zone may quench the He-burning in some models in the core for a while or at least significantly reduce the He-burning core (a process called CNO shell boost, as in Ekström et al., 2008). This has an impact on the nature of the stellar remnant and on nucleosynthesis.
@@
    Models with a mass equal to or larger than 60 $M_{\odot}$ spend more than half of their MS lifetime at the critical velocity limit. However, the mass lost mechanically remains very modest; the most massive model (120 $M_{\odot}$) loses 2% of its initial mass at the end of its evolution.
@@
    With the physics used here, models starting with 70% of the critical velocity on the ZAMS show no evidence of following a chemically homogeneous evolution trend. Actually, their tracks during the MS phase follow the classical redward evolution with the models rotating faster, showing a greater extension towards lower effective temperatures than slower rotators. The impact of fast rotation on the ionisation outputs appears to be modest in the present models.
@@
    Depending on the initial rotation at the ZAMS, a Pop III star can end its stellar lifetime as a BSG or an RSG. The rapidly rotating models will be RSGs, as opposed to the BSGs from the slower-rotating models (for $M_{ini}>20$ $M_{\odot}$).
@@
This work illustrates the fact that the physics of rotational mixing is a key point that needs to be resolved in order to predict stellar yields, especially for the first generations of stars. The physics of rotation is in our view as important as the impact of close binary evolution. Indeed, if mass transfer in close binaries can strongly impact the evolution of stars, primary nitrogen production is essentially a question related to the internal mixing. Rotation appears to be an excellent candidate for what is driving this mixing (we note that close binary evolution and rotation are tightly intertwined through the exchange of angular momentum between orbits and spins of stars). An interesting feature of the rotational mixing is that it shows a dependence on the metallicity, which is quite in line with the fact that important sources of primary nitrogen are needed only at low metallicities (see e.g. the discussion in Chiappini et al., 2005b).
@@
This work also underlines the importance of the initial rotation of Pop III stars. Moreover, even if we had a good knowledge of the rotational velocity distribution in the first Pop III stars, significant uncertainties persist regarding the physics of rotation. Similar work to that discussed here should be done with other approaches to describe the transport of angular momentum and of chemical species by rotation and also by other types of instabilities (improvement in the physics of convection, internal waves, and magnetic instabilities). There is some hope that testing stellar physics on stars that we can observe in the present-day Universe will allow us to constrain these mechanisms in a still-tighter way. Then one will be able to assume that this physics also applies to Pop III stellar models. At the moment, and likely for a very long time to come, there is no way of directly measuring the rotation periods of Pop III stars from observations. Thus, the only way to get information about that is to observe the consequences of stellar models computed with various initial rotations, ideally with different physics for the transport mechanisms, on the early chemical evolution of galaxies (as the origin of the chemical composition of CEMP stars and of the bulk of the galactic halo stars). A possible exception could be provided by cases of extreme flux amplification due to strong gravitational lensing (e.g. Zackrisson et al., 2015). Recently, such extreme lensing has been invoked to explain the Earendel stellar source (Welch et al., 2022), which has a non-negligible probability of being Pop III, although it is more likely to be a metal-poor, massive Pop II star (Schauer et al., 2022). Another channel for getting information on the evolution of the first massive star generations is through predictions of the rate of various types of SNe (see the reviews by Frebel & Norris, 2015; Meynet & Maeder, 2017, and references therein). Recently, nitrogen-rich high-redshift galaxies have been observed (Oesch et al., 2016). The implications of the present rapidly rotating models in this context are discussed in another paper (Nandal et al., in prep.).
@@ Appendix A Additional figures
A.1 Case of the rapidly rotating 120 $M_{\odot}$ model
Figure 13 compares the variation of $U_{\rm r}$, $\Omega$, $D_{\rm shear}$ and $D_{\rm eff}$ as a function of the Lagrangian mass during the core H-burning phase when the mass fraction of hydrogen $X_{\rm c}$ is equal to 0.35 for the 120 $M_{\odot}$ models with a moderate (red curves) and a fast (green curves) initial rotation. The rapidly rotating model shows stronger meridional currents (see the green curve in the upper left panel), which transport angular momentum from inside to outside. The value of $D_{\rm eff}$ is larger by many orders of magnitudes than $D_{\rm shear}$ in the radiative envelope and governs the mixing of the elements in that region.
Figure 13: Variation of $U_{\rm r}$ (upper left panel), $\Omega$ (upper right panel), $D_{\rm shear}$ (solid curves) and $D_{\rm eff}$ (dashed curves) (lower panel) as a function of the Lagrangian mass coordinate in the Pop III 120 $M_{\odot}$ models with different initial rotations. The moderately rotating model was taken from Murphy et al. (2021a).
@@ A.2 CNO shell boost in a Pop III 60 $M_{\odot}$ model
Figure 14: Comparing stages for the 60 $M_{\odot}$ rapidly rotating model as in Fig. 4.
The upper panel of Fig. 14 shows the HRD of the rapidly rotating Pop III 60 $M_{\odot}$ model. Coloured points are overplotted. The orange point indicates when the core He-burning phase begins, while the purple point indicated the stage when an extended intermediate convective zone appears as a result of injections of carbon and oxygen from the He-burning core to the H-burning shell. Meanwhile, the convective He-burning core is reduced. Those points correspond to the vertical coloured lines in the lower Kippenhahn plot, respectively.
@@
In Fig. 15, we have an example of 20 $M_{\odot}$ Pop III models with zero and moderate rotation. We notice that the H- and He-burning shells are very close to each other in the non-rotating model. More precisely, in the non-rotating model, the mass separating the H- to the He-burning shell is a fraction of 1 $M_{\odot}$ during the core C-burning phase. While, in the moderately rotating model, during the same phase, the mass separating the H- and He-burning shell is more than 2 $M_{\odot}$. This is in our view an indirect consequence of the fact that in the moderately rotating model (right panel), a small convective zone occurs above the H-shell at the very beginning of the core He-burning phase (due to the rotational mixing some C and O migrate into the H-shell and boost its energy production). This has for an effect to make the mass of the convective He-core smaller during the first part of the He-burning phase. If the H-shell is replenished in hydrogen by the convective shell, it stays longer in deeper regions of the star, so it migrates slowly towards outer regions. In the non-rotating model, the H-burning shell at the beginning of the core He-burning phase migrates much faster outwards than in the rotating model (the shell is not replenished by the intermediate convective zone and thus progresses more rapidly outwards). This makes the He-core to also increases more rapidly in mass. As a result of this behaviour the two shells, just after the end of the core He-burning phase are separated by less mass in the non-rotating model than in the rotating one. Compare, for instance, in the two Kippenhahn diagram (see Fig. 15) the mass between the blue and green solid lines at a time 4. Due to the greater proximity in the non-rotating model, a CNO shell boost occurs. As a result, we see that the H-shell shifts to a position at a smaller Lagrangian mass coordinate. So we see how complex can be this behaviour. In the rotating model a modest early CNO boost occurs that create conditions that disfavour any further CNO shell boost. In the rotating model the early CNO shell boost does not occur in absence of any rotational mixing, but it creates conditions more favourable for a CNO boost occurring at the end of the core He-burning phase.
Moreover, at the end of core He-burning phase, the moderately rotating mode has higher nitrogen yield (approximately the moderately rotating has ${}^{14}{\rm N}=3.2\times 10^{-7}$, while the non-rotating has ${}^{14}{\rm N}=1.2\times 10^{-7}$). Therefore, we see that those yields on the Fig. 9 are a result after the core He-burning phase.